The sound from a mosquito is produced when it vibrates its wings at an average rate of 500 vibrations per second. What is the time period of the vibration ?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Time taken for 500 vibrations = 1 second Time taken for 1 vibration = 1/500 second. ∴ Time period = 1/500 second.